OpenStudy (anonymous):

If tanθ =1/3 , then secθ = _____.

Nnesha (nnesha):

\[\rm sin = \frac{ opposite }{ hypotenuse }~~~~ \cos = \frac{ adjacent }{ hypotenuse } ~~\tan= \frac{ opposite }{ adjacent }\]

Nnesha (nnesha):



Nnesha (nnesha):

now tan = opposite /adjacent so \[\huge\rm \tan \theta=\frac{ opposite }{ adjacent } =\frac{ 1 }{ 3}\]

Nnesha (nnesha):

 now apply Pythagorean theorem to find 3rd side of the triangle \[\huge\rm a^2+b^2=c^2\] let me know what you get